A gallon of paint will cover 300 sq ft. what lenght of an 8 ft high wall can be covered with 1 gallon of paint?
Answer by edjones(7311)

8x=300
x=300/8=37.5 ft Divide each side by 8.
check:
8*37.5=300
Ed